# Codenames as a Benchmark for Large Language Models

### v1

@@ V-B Two Teams Version
  • Assassin losses: Percentage of games lost due to selecting the assassin word, for the red and blue teams (respectively).
@@ VI-A1 Single Team Version
Looking first at the performance of the LLMs for the single team version of Codenames, see Table I, we can see that the o1-preview model achieved the best performance when paired with a matching model (i.e., the same base LLM for both the codemaster and guesser). The statistical significance of this result was confirmed using a one-sided t-test ($p<0.05$), compared pairwise against all other models. Looking at the other LLMs, the difference in performance is not as clear, with the general trend being that model pairs with a higher loss percentage typically received a higher average score. This can largely be attributed to fact that a loss results in a very high score of 25, meaning that it is often strategically better to play it safe rather than running the risk of selecting the assassin word (at least when averaging performance over multiple trials).
@@
Looking at the other statistics for each LLM, we can make several observations about the different play styles of each model. One major differentiating factor between the models was their level of risk taking as both codemaster and guesser. The riskiness of codemasters was determined using the Clues Avg column. Based on the average clue number given by each codemaster, we can order the LLMs from most to least risky (o1-mini, Sonnet-3.5, o1-preview, Gemini-1.5, GPT-4o, and Llama-3.1). This riskiness in terms of the clue number given, also heavily correlates with the percentage number of losses (i.e., riskier codemasters tended to have a high percentage number of losses, when accounting for differences in the guesser model). The Median and Min columns also back up these findings as these values appear generally lower for the risky codemasters, indicating that when these models do win that they do so in a lower number of turns. Risky codemasters are also likely to have a higher number of Blue and Civillian words selected.
@@
The riskiness of guessers was estimations using a combination of the Stop Early and Stop Late values for each model (averaged across the different possible codemasters). A higher Stop Early percentage should indicate a more cautious guesser, while a higher Stop Late percentage would indicate a more risky guesser. Note this way of measuring guesser riskiness is not an exact science, as it also appears that some models (such as GPT-4o) are much more likely to stick to the clue number provided by the codemaster. Based on the general trends of these metrics, it would appear that Gemini-1.5 and Llama-3.1 are more likely to stop early. Models also appeared to have a higher Stop Early percentage when paired with a risky codemaster (e.g., Sonnet-3.5) and were much lower for cautious codemasters (e.g., Llama-3.1). The o1-preview and o1-mini LLMs had a very high stop early percentage, although this could have been due to the fact that they were only evaluated when paired with themselves and are also risky codemasters. However, both o1-preview and o1-mini also had very high stop late percentages, indicating that these models are less likely to stick to the codemaster’s clue number. Sonnet-3.5 and Llama-3.1 did appear to have a slightly higher stop Late percentage than other LLMs, when accounting for different codemasters, but this was still fairly low by comparison.
@@
In contrast to the LLM agents, the word-vector agents consistently achieved better performance when paired with a matching partner model, resulting in a very low mean score and a 0% loss rate. GloVe appeared to have a slightly higher average clue number than the other models, which also resulted in a slightly lower mean score. However, it is apparent that the performance of these models drops significantly when playing with a different word-vector approach (particularly when Word2Vec is paired with GloVe). This limitation does not appear to be present in the LLM agents, which performed equally well when paired with other LLMs.
@@ VI-A2 Two Teams Version
Looking at the result for the two teams version of Codenames, see Table II, we can see some interesting differences in the performance of certain LLMs compared to the single team version. Firstly, in would appear that riskier codemasters (e.g., sonnet-3.5) perform much better, likely due to the reduced impact that selecting an assassin word has on a team’s average win-rate. Looking at the overall win-rates for both the red and blue teams, there does not appear to be a substantial difference between each side (the red team always goes first, but has nine words to identify compared to blue team’s eight words). Across all trials, Sonnet-3.5 had the highest average win-rate for the red team (56.75%), while GPT-4o has the highest overall win-rate for the blue team (58.75%). The percentage of assassin losses for each model did not appear substantially different from in the single player version, indicating little change in terms of the overall riskiness / cautiousness of each LLM’s strategy. For the word-vector agents, Word2Vec was the worst performing overall while GloVe was the best performing (same result as for the single team version).
@@ VI-B Qualitative Observations
This section will discuss some anecdotal findings regarding model behaviour that was observed throughout the games played, including several instances where it appeared that the LLMs did not fully internalise the game’s rules. We propose some possible explanations for these behaviours, although the black box nature of LLMs makes it difficult to verify the true cause.
@@ VI-B2 Models giving invalid responses
Codemasters giving invalid responses: Invalid responses from LLM codemasters often came in the form of clue words that contained, or were contained within, another word on the board. For example, trying to give the clue Bonfire or Campfire when FIRE is a word on the board would be considered an invalid response. Due to the way that our invalid clue checker was implemented, essentially checking for substring matches between the provided clue and all remaining words on the board, there is a risk for some potentially valid clues to be incorrectly flagged as invalid. For example, the clue Education was flagged as invalid because it contained the board word CAT, even though most human players would argue that the word education is not derived from the word cat and thus doesn’t violate the game’s official rules. One way to address this issue in future would be to employ a more advanced checker, perhaps utilising a database of compound words that derive from the given clue, although this may not completely fix the problem.
@@ VI-B3 Idiosyncrasies of model behaviour
Over emphasis on first word: Firstly, many codemaster agents would overemphasise their clue’s connection to one word on the board, neglecting strong connections with the other words they intended to hint at. For example, the GPT-4o model once gave the clue (Picnic, 2) with the remaining team words to identify being PLATE, TRUNK, and BEACH. While the word has strong associations with PLATE, which was correctly guessed, connections to the other words seem tenuous at best. This clue was also given when the incorrect word JAM was on the board, which was picked by the guesser after PLATE. This first word prioritisation by codemasters was a repeatedly observed behaviour across all LLMs.
@@ VII Future work
If future, if LLM agents can be developed with a high level of competence Codenames, we may also be able to extract explainable strategies by studying their gameplay. This approach is already being taken in other AI dominated games such as Go, Chess and Dota 2 [45].
@@ VIII Conclusion
In this paper we have explored the potential of the game Codenames as a suitable benchmark for assessing the language reasoning capabilities of multiple LLM agents, as well as their strategic tendencies within the game. We did this by having different LLMs play both the single team and two team version of Codenames, alongside a variety of different teammates and opponents. We found that each LLM exhibits a unique emergent style of play and that they do not necessarily perform best when paired together. In fact, each LLM often performs better in a specific role (either as codemaster or guesser). We found that a cautious playstyle often resulted in high performance for the single team version, with the exception of the new o1-preview model which was able to achieve superior performance over all other LLMs despite exhibiting a more risky strategy. For the two teams version, playing risky often led to a higher average win-rate against a more cautious opponent, suggesting that being slightly more risky than the opposing team is the preferred behaviour. We also discuss why LLM agents are more suitable for human play than previous word-vector approaches, observed idiosyncrasies in model behaviour, and the rich potential for future work in LLM research utilising Codenames.

### v2

@@ V-B Two Teams Version
  • Assassin losses: Percentage of games lost due to selecting the assassin word, for the red and blue teams respectively.
@@ VI-A1 Single Team Version
Looking first at the performance of the LLMs for the single team version of Codenames, see Table I, we can see that the o1-preview model achieved the best performance when paired with a matching model (i.e., the same base LLM for both the codemaster and guesser). Looking at the other LLMs, the difference in performance is not as clear, with the general trend being that model pairs with a higher loss percentage typically received a higher average score. This can largely be attributed to fact that a loss results in a very high score of 25, meaning that it is often strategically better to play it safe rather than running the risk of selecting the assassin word (at least when averaging performance over multiple trials).
@@
Looking at the other statistics for each LLM, we can make several observations about the different play styles of each model. One major differentiating factor between the models was their level of risk taking as both codemaster and guesser. The riskiness of codemasters was estimated using the Clues Avg column. Based on the average clue number given by each codemaster, we can order the LLMs from most to least risky (o3-mini, DeepSeek-R1, o1-mini, Sonnet-3.5, o1-preview, DeepSeek-V3, Gemini-1.5, GPT-4o, and Llama-3.1). This riskiness in terms of the clue number given also heavily correlates with the percentage number of losses, with a Spearman’s rank correlation coefficient of 0.821 between these two columns when considering all LLM model pairs (i.e., riskier codemasters tended to have a high percentage number of losses, when accounting for differences in the guesser model). The Median and Min columns also back up these findings as these values appear generally lower for the risky codemasters, indicating that when these models do win that they do so in a lower number of turns. Risky codemasters are also likely to have a higher number of Blue and Civillian words selected.
@@
The riskiness of guessers was estimated using a combination of the Stop Early and Stop Late values for each model (averaged across the different possible codemasters). A higher Stop Early percentage should indicate a more cautious guesser, while a higher Stop Late percentage would indicate a more risky guesser. Note this way of measuring guesser riskiness is not an exact science, as it also appears that some models (such as GPT-4o) are much more likely to stick to the clue number provided by the codemaster. Based on the general trends of these metrics, it would appear that Gemini-1.5 and Llama-3.1 are more likely to stop early. Models also appeared to have a higher Stop Early percentage when paired with a risky codemaster (e.g., Sonnet-3.5) and were much lower for cautious codemasters (e.g., Llama-3.1). The o1-preview, o1-mini, o3-mini, DeepSeek-R1 and DeepSeek-V3 LLMs also had very high Stop Early percentages, although this could have been due to the fact that they were only evaluated when paired with themselves and are also risky codemasters. However, both o1-preview and o1-mini also had very high Stop Late percentages, indicating that these models are less likely to stick to the codemaster’s clue number. Sonnet-3.5 and Llama-3.1 did appear to have a slightly higher Stop Late percentage than other LLMs, when accounting for different codemasters, but this was still fairly low by comparison.
@@
In contrast to the LLM agents, the word-vector agents consistently achieved better performance when paired with a matching partner model, resulting in a very low mean score and a 0% loss rate. GloVe appeared to have a slightly higher average clue number than the other models, which also resulted in a slightly lower mean score. However, it is apparent that the performance of these models drops significantly when playing with a different word-vector approach (particularly when Word2Vec is paired with GloVe). This limitation does not appear to be present in the LLM agents, which performed roughly equally well when paired with other LLMs.
@@ VI-A2 Two Teams Version
Looking at the result for the two teams version of Codenames, see Table II, we can see some interesting differences in the performance of certain LLMs compared to the single team version. Firstly, it would appear that riskier codemasters (e.g., sonnet-3.5) perform much better, likely due to the reduced impact that selecting an assassin word has on a team’s average win-rate. Looking at the overall win-rates for both the red and blue teams, there does not appear to be a substantial difference between each side (the red team always goes first, but has nine words to identify compared to blue team’s eight words). Across all trials, Sonnet-3.5 had the highest average win-rate for the red team of 56.75% (compared to 51.00% for GPT-4o, 41.50% for Gemini-1.5, and 46.25% for Llama-3.1) while GPT-4o has the highest overall win-rate for the blue team of 58.75% (compared to 46.25% for Gemini-1.5, 50.25% for Sonnet-3.5, and 49.25% for Llama-3.1). The percentage of assassin losses for each model did not appear substantially different from in the single team version, indicating little change in terms of the overall riskiness / cautiousness of each LLM’s strategy. For the word-vector agents, Word2Vec was the worst performing overall while GloVe was the best performing (same result as for the single team version).
The observed difference in the effectiveness of different models / playstyles between the single and two teams versions of Codenames demonstrates some of the potential limitations with relying on the scoring approach of the single team version. Most notably, selecting the assassin word has a huge impact on a team’s overall mean score for the single team version, which results in a game that encourages cautious play. The two team version of Codenames instead appears to incentivise the opposite, with Sonnet-3.5 going from the worst performing LLM for the single team version (in terms of mean score across all games played) to the best performing LLM for the two teams version (based on average win-rate across all games played). This would seem to imply that playing slightly riskier than your opponent is a smart strategy for this version, particularly when paired against overly cautious teams.
@@ VI-B Qualitative Observations
This section will discuss anecdotal findings regarding model behaviour that was observed throughout the games played, including several instances where it appeared the LLMs did not fully internalise the game’s rules. We propose some possible explanations for these behaviours, although the black box nature of LLMs makes it difficult to verify the true cause.
@@ VI-B2 Models giving invalid responses
Codemasters giving invalid responses: Invalid responses from LLM codemasters often came in the form of clue words that contained, or were contained within, another word on the board. For example, trying to give the clue Bonfire or Campfire when FIRE is a word on the board would be considered an invalid response. Due to the way that our invalid clue checker was implemented, essentially checking for substring matches between the provided clue and all remaining words on the board, there is a risk for some potentially valid clues to be incorrectly flagged as invalid. For example, the clue Education was flagged as invalid because it contained the board word CAT, even though most human players would argue that the word education is not derived from the word cat and thus doesn’t violate the game’s official rules. One way to address this issue in the future would be to employ a more advanced checker, perhaps utilising a database of compound words that derive from the given clue, although this may not completely fix the problem.
@@ VI-B3 Idiosyncrasies of model behaviour
Over emphasis on one word: Firstly, many codemaster agents would overemphasise their clue’s connection to one word on the board, neglecting strong connections with the other words they intended to hint at. For example, the GPT-4o model once gave the clue (Picnic, 2) with the remaining team words to identify being PLATE, TRUNK, and BEACH. While the word has strong associations with PLATE, which was correctly guessed, connections to the other words seem tenuous at best. This clue was also given when the incorrect word JAM was on the board, which was picked by the guesser after PLATE. This single word prioritisation by codemasters was a repeatedly observed behaviour across all LLMs.
@@ VII Future work
Additional experiments could provide further comparisons on how different techniques (including both LLM and word-vector approaches) can be most effectively combined. For example, while the current word-vector agents cannot reliably act as guesser when paired with an LLM, due to their inability to interpret words outside of their provided corpus, they can still operate as Codemaster alongside an LLM guesser. Such a comparison could help to indicate if any of the word-vector approaches provide clues that are more human interpretable, as LLMs have previously been shown to work well at collaborating alongside human players [28].
@@
In addition to Codenames, many other language-based boardgames could also be explored in future studies to help verify if our LLM performance and behavioural findings are specific only to codenames or are more general. Some example games that could be investigated further include “So Clover”, ‘Just One”, “Letter Jam”, “Decrypto”, “Medium” and “Master Word”. If LLM agents can be developed with a high level of competence in these language-based games, we may also be able to extract explainable strategies by studying their gameplay. This approach is already being taken in other AI dominated games such as Go, Chess and Dota 2 [45].
@@ VIII Conclusion
In this paper we have explored the potential of the game of Codenames as a suitable benchmark for assessing the language reasoning capabilities of multiple LLM agents, as well as their strategic tendencies within the game. We did this by having different LLMs play both the single team and two team version of Codenames, alongside a variety of different teammates and opponents. We found that each LLM exhibits a unique emergent style of play and that they do not necessarily perform best when paired together. In fact, each LLM often performs better in a specific role (either as codemaster or guesser). We found that a cautious playstyle often resulted in high performance for the single team version, with the exception of the new o1-preview model which was able to achieve superior performance over all other LLMs despite exhibiting a more risky strategy. For the two teams version the opposite appeared to be the case, where playing risky often led to a higher average win-rate against a more cautious opponent. We also discuss why LLM agents are more suitable for human play than previous word-vector approaches, observed idiosyncrasies in model behaviour, and the rich potential for future work in LLM research utilising Codenames.